what is the solution to the system shown below?
2x+y=-3
x-5y=4
A. (0,-3)
B. (-3,4)
C. (-1,-1)
D. (,-5)

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

what is the solution to the system shown below?
2x+y=-3
x-5y=4
--------------
Multiply the 2nd equation by 2, then subtract.

Answer by rothauserc(4718)

1) 2x +y = -3
2) x -5y = 4
:
solve equation 2 for x
:
x = 5y +4
:
substitute for x in equation 1
:
2(5y +4) + y = -3
:
10y +8 +y = -3
:
11y = -11
:
y = -1
x = 5(-1) +4 = -5 +4 = -1
:
***************************************
solution is (-1, -1) and answer is C
:check our answer with equations 1 and 2
:
1) 2(-1) +(-1) = -3
-2 -1 = -3
-3 = -3
:
2) -1 -5(-1) = 4
-1 +5 = 4
4 = 4
:
our answer checks :-)
****************************************
: